# PBPK-iPINNs: Inverse Physics-Informed Neural Networks for Physiologically Based Pharmacokinetic Brain Models

### v2

@@ 6 Conclusion
To the best of our knowledge, research on inverse physics-informed neural networks (iPINNs) applied to physiologically based pharmacokinetic (PBPK) modeling is limited, and no study has applied iPINNs to the 4-compartment, permeability-limited PBPK brain model used in this work for estimating drug- and patient-specific parameters and predicting drug concentration profiles. Experimentally determining drug-specific or patient-specific parameters can be a complex and time-consuming process due to challenges in sampling and limitations of analytical technologies (in vitro experiments). Existing numerical and statistical algorithms often fail to converge when the system of ODEs becomes highly stiff.

### v3

@@ 6 Conclusion
To the best of our knowledge, research on inverse physics-informed neural networks (iPINNs) applied to physiologically based pharmacokinetic (PBPK) modeling is limited, and no study has applied iPINNs to the 4-compartment, permeability-limited PBPK brain model used in this work for estimating drug- and patient-specific parameters and predicting drug concentration profiles.
@@
Limitations and Future Work
Experimentally determining drug-specific or patient-specific parameters can be a complex and time-consuming process due to challenges in sampling and limitations of analytical technologies (e.g., in vitro experiments). Furthermore, existing numerical and statistical algorithms often fail to converge when the system of ODEs becomes highly stiff. Since this study recommends PBPK-iPINN as an alternative to conventional numerical and statistical methods, we plan to apply the same methodology described in this paper to metabolic flux analysis (MFA) in future work. MFA consists of a large number of differential equations (more than 60 ODEs) and more than 80 parameters to be estimated, where we have already observed that traditional numerical and statistical methods fail to converge.
@@
Author Contributions
C.W. wrote the manuscript; C.W. designed the research; C.W., K.W., and P.R. performed the PINN simulations; C.W. and N.H. analyzed the data and conducted the parameter identifiability analysis.